If the length of a rectangle is 5 times its width and if its area is 180, what is its perimeter?
Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
(5w)w = 180
5w^2 = 180
w^2 = 36
w = 6 (Tossing out the negative solution for unit measure)
w = 6 and L = 30
P = 2L + 2w = 60 + 12 = 72units